A patio door is 72 inches wide. If a drawing of the door is made with a scale of 2 in. = 3 ft, what is the width of the door on the drawing? (Hint: 12 inches = 1 foot).
Answer by tenkun(12)

12 inches = 1 foot
72 inches = 72/12 =6 ft
3 ft =2 in
6 ft= 2/3* 6
=4 in